# Dark gauge-mediated supersymmetry breaking with a massless dark photon

### v1

@@ 2 Supersymmetric kinetic mixing
where $\hat{X}$ and $\hat{B}$ are the physical states after the transformation. While any choice of $\omega$ can diagonalize the kinetic term, we adopt the following basis Holdom:1985ag ; Fabbrichesi:2020wbt for our analysis:
@@
This transformation leads to the diagonalized states of the system, effectively illustrating how kinetic mixing impacts the physical properties of the gauge bosons. In this basis, the gauge boson $B$ can be interpreted as the physical $U(1)_{Y}$ gauge boson since SM particles interact exclusively with $B$.
In this supersymmetric framework, component fields such as gauge bosons, gauginos, and auxiliary fields can be coherently rotated by a unified basis transformation. We note that one can take different rotation angle $\omega$ for each component, without changing any physical results. However, this alignment facilitates a clear delineation of interaction terms in the Lagrangian, even after diagonalizing the kinetic mixing terms. The interaction terms for the gauge boson in this basis are expressed as
@@
where $g_{Y}=g_{Y}^{\prime}/\sqrt{1-\epsilon^{2}}$ represents the effective $U(1)_{Y}$ coupling after the diagonalization, and $g_{D}$ is $U(1)_{D}$ coupling. $Y$ denotes the hypercharge associated with the visible current $J^{\mu}_{Y}$, and $D$ represents the $U(1)_{D}$ charge of the dark current $J^{\mu}_{D}$. If $J_{Y}^{\mu}$ and $J^{\mu}_{D}$ take the same form, the hypercharge is effectively modified by the contribution from the dark gauge symmetry as
@@
This suggests that within the GMSB framework, messengers that carry dark charge could influence the generation of mass for visible superpartners. A quantitative analysis will be provided in Section 4. In general, Eq. (7) indicates that any particles with a non-zero dark charge will have a fractional hypercharge Holdom:1985ag ; Holdom:1986eq . As a result, constraints on a massless dark photon primarily arise from the presence of milli-charged exotic particles Prinz:1998ua ; Davidson:2000hf ; PhysRevD.75.032004 ; Jaeckel:2012yz ; Vogel:2013raa ; Chang:2018rso ; Magill:2018tbb ; Stebbins:2019xjr . Therefore, we assume that there are no light states with a non-zero dark charge, allowing the kinetic mixing parameter $\epsilon$ to remain largely unconstrained under these constraints.
@@
Kinetic mixing can affect electroweak symmetry breaking (EWSB), as the value of the $\mu$ parameter is determined by a combination of Higgs mass parameters to ensure EWSB. So, at large $\epsilon$ values, the Higgs mass parameter at the low-energy scale may not satisfy the EWSB condition. This detail is discussed in Section 4 and Appendix A. These aspects will be further explored in subsequent sections to assess the impact of large $\epsilon$. Also, caution is necessary when considering large values of $\epsilon$ and $g_{D}$ from the perspective of UV physics. Elevated coupling strengths could lead to the emergence of Landau poles at energy scales near those of IR physics, an issue addressed in Appendix B. Consequently, kinetic mixing cannot be arbitrarily large, although we demonstrate that substantial kinetic mixing remains viable.
@@ 4.1 Higgs mass and electroweak symmetry breaking
The $\mu$-term is a supersymmetric mass parameter appearing in the superpotential that couples the two Higgs doublets, while the $b_{\mu}$-term is a soft supersymmetry-breaking bilinear term in the Higgs potential. In the conventional GMSB, EWSB can occur through the renormalization group (RG) running of $m_{H_{u}}^{2}$ and $m_{H_{d}}^{2}$, which generates the hierarchy $m_{H_{u}}^{2}<m_{H_{d}}^{2}$ at the EW scale. This is the case in dark GMSB as well. [^1] In both conventional GMSB and dark GMSB, $m_{H_{u}}^{2}=m_{H_{d}}^{2}$ at the messenger scale, and thus cannot satisfy Eq. (27). A key difference in dark GMSB, however, is that the bino and mixed bino-dark photino soft masses can contribute significantly to the running of $m_{H_{u}}^{2}$ and $m_{H_{d}}^{2}$. Large values of $g_{D}$ and $\epsilon$ significantly enhance these soft masses parameters, eventually preventing $m_{H_{u}}^{2}$ from being driven to negative values. (See Eqs. (38), (39), and Appendix A.) This implies that the RHS of Eq. (27) becomes negative for large $g_{D}$ and $\epsilon$, and therefore must be constrained. On the other hand, this also implies that $|\mu|$ could be significantly smaller than in conventional GMSB for moderate values of $g_{D}$ and $\epsilon$, which brings down the mass of the higgsino.
Figure 3: (a) Isocontours of $|\mu|$ in the $\tan\beta-\epsilon$ plane for a complete $SU(5)$ messenger multiplet (Scenario I). Here, $F/M$ is adjusted to consistently achieve $m_{h^{0}}=125$ GeV, with $F/M_{\mathrm{mess}}^{2}=2/3$ fixed. The EW vacuum cannot be achieved in the gray region, which is therefore ruled out. (b) The variation of $|\mu|$ with changes in $\epsilon$. This shows that $\mu$ sharply turns imaginary as $\epsilon$ increases, obstructing successful EWSB. We take $F/M_{\mathrm{mess}}=800$ TeV, $M_{\mathrm{mess}}=1200$ TeV, $g_{D}=0.4$, and $\tan\beta=15$, values compatible with our benchmark point (see Table 2).
@@
Figure 3 illustrates the effect of dark GMSB on $|\mu|$ in detail. Figure 3(a) shows the values of $|\mu|$ as a function of $\tan\beta$ and $\epsilon$ needed to achieve EWSB and the correct Higgs mass. Since $1/\cos(2\beta)$ diverges when $\tan\beta\rightarrow 1$, the contour of $|\mu|$ becomes parallel to the constant $\tan\beta$ line. This figure also shows one crucial feature of dark GMSB, which is that $|\mu|$ is affected by $\epsilon$. The value of $|\mu|$ decreases as $\epsilon$ increases, with the variation becoming more sensitive at larger $\epsilon$, as shown in Figure 3(b). The boundary of the gray region is where $|\mu|\simeq 0$. If the values of $g_{D}$ and $\epsilon$ lie closer to the gray region, it makes $|\mu|$ to be small. Therefore, the spectrum of dark GMSB can contain a light higgsino, with mass determined by $|\mu|$, which will be demonstrated in Section 5.
@@
The dark gauge mediation effect can induce a small fractional change in the Higgs mass through the change of soft stop mass, Eqs. (20, 21). The effect can be explicitly written by
@@ 6 Particle spectra and implications for phenomenology
Figure 16 shows isocontours of the branching ratio of the Higgs to the lightest neutralinos for negative and positive $\epsilon$ (left and right panels, respectively). A sharp rise in the branching ratio is observed near the boundary of the EW vacuum constraint, where $\mu$ decreases rapidly and the Higgsino components of $\tilde{N}^{0}$ correspondingly increase. Furthermore, it is evident that the branching ratio increases as $\tan\beta$ becomes smaller. This is because the NLSP contains a relatively larger $\tilde{H}_{d}$ component, and since $\sin\alpha\simeq\cos\beta\simeq 1/\tan\beta$ in the $m_{H^{0}}\gg m_{h^{0}}$ and large $\tan\beta$ limits, the $N_{0\tilde{H}_{d}}\sin\alpha$ term Eq. (49) grows as $\tan\beta$ decreases. Interestingly, some of the parameter space of $\tan\beta$ and $\epsilon$ is accessible to the future Higgs factories such as FCC-ee FCC:2018evy , ILC ILC:2007oiw , or CEPC CEPCStudyGroup:2023quu where we have taken $\text{Br}_{\text{inv}}\simeq 0.19\%$ as a projected upper bound on the invisible Higgs branching ratio deBlas:2019rxi .
@@ 7.2 Relic Messengers
However, the relic density of the messengers could be substantially suppressed if the reheating temperature $T_{\mathrm{reh}}$ is significantly lower than $m_{\Psi}$. Then, the production of the messengers is inhibited by the large Boltzmann factor, $e^{-m_{\Psi}/T_{\mathrm{max}}}$. [^5] The maximum temperature during reheating, $T_{\mathrm{max}}$, can be much higher than $T_{\mathrm{reh}}$, which is the temperature at the end of reheating. Once the temperature during reheating reaches $T_{\mathrm{max}}$, it decreases proportionally to $T\propto T_{\mathrm{max}}a^{-3/8}$, where $a$ is the scale factor. Even if the messengers achieve thermal equilibrium with the thermal bath, they would decouple during reheating, and their number density would be diluted by subsequent entropy production Giudice:2000ex . If $T_{\mathrm{max}}\simeq m_{\Psi}$, the relic density of the messengers is given by Giudice:2000ex
@@ Appendix B Evolution gauge couplings and kinetic mixing at high scales
In this appendix, we discuss the evolution of the gauge coupling constants and kinetic mixing in the dark GMSB model above the messenger scale. In the original basis of Eqs. (1,2), the gauge interaction of the matter fields can be expressed as
@@
where the trace is taken over all chiral multiplets.
Figure 19: The region of parameter space in which $g_{D}$ or $\epsilon$ exhibits a Landau pole below $Q=100M_{\mathrm{mess}}$ (gray shaded). The solid curve represents the bound for Scenario I, while the dashed curve indicates the bound for Scenario II. A large kinetic mixing, $\epsilon\lesssim 0.7$, is allowed for $g_{D}\simeq 1$. Also shown are isocontours of the effective coupling $g_{\mathrm{eff}}=g_{D}\epsilon/\sqrt{(1-\epsilon^{2})}$.

### v2

@@ 2 Supersymmetric kinetic mixing
where $\hat{X}$ and $\hat{B}$ are the physical states after the transformation. [^1] A requirement that the kinetic terms of the gauge bosons be positive constrains the kinetic mixing to be less than 1 Burgess:2008ri ; Gan:2023jbs . While any choice of $\omega$ can diagonalize the kinetic term and is physically equivalent Pan:2018dmu , we adopt the following basis Holdom:1985ag ; Fabbrichesi:2020wbt for our analysis: [^2] Although renormalization group running may change the value of $\epsilon$, one can always take the basis of Eq. (4) provided that $U(1)_{D}$ remains unbroken.
@@
This transformation leads to the diagonalized states of the system, effectively illustrating how kinetic mixing impacts the physical properties of the gauge bosons. In this supersymmetric framework, component fields such as gauge bosons, gauginos, and auxiliary fields can be coherently rotated by a unified basis transformation. We note that one can take different rotation angle $\omega$ for each component, without changing any physical results. However, this alignment facilitates a clear delineation of interaction terms in the Lagrangian, even after diagonalizing the kinetic mixing terms.
In this basis, the gauge boson $B$ can be interpreted as the physical $U(1)_{Y}$ gauge boson since SM particles interact exclusively with $B$. The interaction terms for the gauge boson in this basis are expressed as
@@
where $g_{Y}=g_{Y}^{\prime}/\sqrt{1-\epsilon^{2}}$ represents the effective $U(1)_{Y}$ coupling after the diagonalization, and $g_{D}$ is $U(1)_{D}$ coupling. $Y$ denotes the hypercharge associated with the visible current $J^{\mu}_{Y}$, and $D$ represents the $U(1)_{D}$ charge of the dark current $J^{\mu}_{D}$. Therefore, this basis isolates the specific linear combination of the gauge fields that is invisible to SM particles as the $X$ field, simplifying the interpretation of experimental results.
If $J_{Y}^{\mu}$ and $J^{\mu}_{D}$ take the same form, the hypercharge is effectively modified by the contribution from the dark gauge symmetry as
@@
This suggests that within the GMSB framework, messengers that carry dark charge could influence the generation of mass for visible superpartners. A quantitative analysis will be provided in Section 4. In general, Eq. (7) indicates that any particles with a non-zero dark charge will have a fractional hypercharge Holdom:1985ag ; Holdom:1986eq . As a result, constraints on a massless dark photon primarily arise from the presence of milli-charged exotic particles Prinz:1998ua ; Davidson:2000hf ; PhysRevD.75.032004 ; Jaeckel:2012yz ; Vogel:2013raa ; Chang:2018rso ; Magill:2018tbb ; Stebbins:2019xjr ; Fiorillo:2024upk . Therefore, we assume that there are no light states with a non-zero dark charge, allowing the kinetic mixing parameter $\epsilon$ to remain largely unconstrained under these constraints.
@@
Kinetic mixing can affect electroweak symmetry breaking (EWSB), as the value of the $\mu$ parameter is determined by a combination of Higgs mass parameters to ensure EWSB. So, at large $\epsilon$ values, the Higgs mass parameter at the low-energy scale may not satisfy the EWSB condition. This detail is discussed in Section 4 and Appendix A. These aspects will be further explored in subsequent sections to assess the impact of large $\epsilon$. Caution is necessary when considering large values of $\epsilon$ and $g_{D}$, and, in particular, the couplings may diverge in the $\epsilon\rightarrow 1$ limit (see Eq. (7)) leading to a non-perturbative regime and a breakdown of perturbative unitarity. Moreover, from the perspective of UV physics, elevated coupling strengths could lead to the emergence of Landau poles at energy scales close to those of IR physics, an issue addressed in Appendix B. Consequently, kinetic mixing cannot be arbitrarily large, although we demonstrate that substantial kinetic mixing remains viable.
@@ 4.1 Higgs mass and electroweak symmetry breaking
The $\mu$-term is a supersymmetric mass parameter appearing in the superpotential that couples the two Higgs doublets, while the $b_{\mu}$-term is a soft supersymmetry-breaking bilinear term in the Higgs potential. Generating $\mu$ and $b_{\mu}$ terms required by the above relations for EWSB without fine tuning is an acute problem in the GMSB Giudice:1998bp . In this paper, we do not address this issue directly; instead, as commonly done in typical studies of GMSB, we assume that $\mu$ and $b_{\mu}$ can be generated to satisfy the relations in Eqs. (26) and (27). In the conventional GMSB, EWSB can occur through the renormalization group (RG) running of $m_{H_{u}}^{2}$ and $m_{H_{d}}^{2}$, which generates the hierarchy $m_{H_{u}}^{2}<m_{H_{d}}^{2}$ at the EW scale. This is the case in dark GMSB as well. [^3] In both conventional GMSB and dark GMSB, $m_{H_{u}}^{2}=m_{H_{d}}^{2}$ at the messenger scale, and thus cannot satisfy Eq. (27). A key difference in dark GMSB, however, is that the bino and mixed bino-dark photino soft masses can contribute significantly to the running of $m_{H_{u}}^{2}$ and $m_{H_{d}}^{2}$. Large values of $g_{D}$ and $\epsilon$ significantly enhance these soft masses parameters, eventually preventing $m_{H_{u}}^{2}$ from being driven to negative values. (See Eqs. (38), (39), and Appendix A.) The one-loop beta functions of $m_{H_{u}}^{2}$ and $m_{H_{d}}^{2}$ receive identical negative contributions from $g_{D}$ and $\epsilon$. Thus, regardless of the value of $\beta$, large $g_{D}$ and $\epsilon$ drive the right-handed side of Eq. (27) to become negative and must therefore be constrained. On the other hand, this also implies that $|\mu|$ could be significantly smaller than in conventional GMSB for moderate values of $g_{D}$ and $\epsilon$, which brings down the mass of the higgsino.
Figure 3: (a) Isocontours of $|\mu|$ in the $\tan\beta-\epsilon$ plane for a complete $SU(5)$ messenger multiplet (Scenario I). Here, $F/M$ is adjusted to consistently achieve $m_{h^{0}}=125$ GeV, with $F/M_{\mathrm{mess}}^{2}=2/3$ fixed. The EW vacuum cannot be achieved in the gray region, which is therefore ruled out. (b) The variation of $|\mu|$ with respect to $\epsilon$ is shown for $g_{D}=0.3$, $0.4$, and $0.5$. This shows that $\mu$ sharply turns imaginary as $\epsilon$ increases, obstructing successful EWSB. We take $F/M_{\mathrm{mess}}=800$ TeV, $M_{\mathrm{mess}}=1200$ TeV, and $\tan\beta=15$ — values that match our benchmark point except for $g_{D}$ (see Table 2).
@@
Figure 3 illustrates the effect of dark GMSB on $|\mu|$ in detail. Figure 3(a) shows the values of $|\mu|$ as a function of $\tan\beta$ and $\epsilon$ needed to achieve EWSB and the correct Higgs mass. Since $1/\cos(2\beta)$ diverges when $\tan\beta\rightarrow 1$, the contour of $|\mu|$ becomes parallel to the constant $\tan\beta$ line. This figure also shows one crucial feature of dark GMSB, which is that $|\mu|$ is affected by $\epsilon$. The value of $|\mu|$ decreases as $\epsilon$ increases, with the variation becoming more sensitive at larger $\epsilon$, as shown in Figure 3(b). The boundary of the gray region is where $|\mu|\simeq 0$. Higher-order loop corrections slightly shift the precise location of the boundary of the gray region; however, the qualitative behavior, where $|\mu|$ steeply decreases near the boundary is expected to remain similar. If the values of $g_{D}$ and $\epsilon$ lie closer to the gray region, it makes $|\mu|$ to be small. Therefore, the spectrum of dark GMSB can contain a light higgsino, with mass determined by $|\mu|$, which will be demonstrated in Section 5.
@@
The dark gauge mediation effect can induce a small fractional change in the Higgs mass through the change of soft stop mass, Eqs. (20), (21). The effect can be explicitly written by
@@ 6 Particle spectra and implications for phenomenology
Figure 16 shows isocontours of the branching ratio of the Higgs to the lightest neutralinos for negative and positive $\epsilon$ (left and right panels, respectively). A sharp rise in the branching ratio is observed near the boundary of the EW vacuum constraint, where $\mu$ decreases rapidly and the Higgsino components of $\tilde{N}^{0}$ correspondingly increase. Furthermore, it is evident that the branching ratio increases as $\tan\beta$ becomes smaller. This is because the NLSP contains a relatively larger $\tilde{H}_{d}$ component, and since $\sin\alpha\simeq\cos\beta\simeq 1/\tan\beta$ in the $m_{H^{0}}\gg m_{h^{0}}$ and large $\tan\beta$ limits, the $N_{0\tilde{H}_{d}}\sin\alpha$ term Eq. (49) grows as $\tan\beta$ decreases. For small $\tan\beta$, a large stop mass, i.e., large $F/M$, is required to achieve the observed Higgs mass of $125$ GeV. This also sufficiently raises the $N_{0}$ mass as $\tan\beta$ becomes smaller than about 5-6 so that the exotic Higgs decays may not occur because $m_{\tilde{N}_{0}}>m_{h_{0}}/2$. Furthermore, the numerical evaluation of the spectrum becomes unstable due to convergence issues below $\tan\beta\lesssim 7$. Interestingly, some of the parameter space of $\tan\beta$ and $\epsilon$ is accessible to the future Higgs factories such as FCC-ee FCC:2018evy , ILC ILC:2007oiw , or CEPC CEPCStudyGroup:2023quu where we have taken $\text{Br}_{\text{inv}}\simeq 0.19\%$ from $\Gamma_{h^{0}}^{\mathrm{tot}}=4.07\times 10^{-3}\,\mathrm{GeV}$ as a projected upper bound on the invisible Higgs branching ratio deBlas:2019rxi .
The invisible decay of $Z$ boson to the $\tilde{N}_{0}$ is also viable with the following decay rate Dreiner:2012ex
$\Gamma(Z\rightarrow\tilde{N}_{0}\tilde{N}_{0})=\frac{g_{2}^{2}m_{Z}}{12\pi}\left(1-\frac{4m_{\tilde{N}_{0}}^{2}}{m_{Z}^{2}}\right)^{3/2}\,\left|\frac{1}{2\cos\theta_{W}\sin\theta_{W}}\left(N_{0\tilde{H}_{u}}^{2}-N_{0\tilde{H}_{d}}^{2}\right)\right|^{2}$ (50)
However, due to the suppressed Higgsino composition in $\tilde{N}_{0}$, the LEP constraint $\Gamma(Z\rightarrow\tilde{N}_{0}\tilde{N}_{0})<3\,\mathrm{MeV}$ from the $Z$ boson invisible decay width ALEPH:2005ab is not relevant.
@@ 7.2 Relic Messengers
However, the relic density of the messengers could be substantially suppressed if the reheating temperature $T_{\mathrm{reh}}$ is significantly lower than $m_{\Psi}$. Then, the production of the messengers is inhibited by the large Boltzmann factor, $e^{-m_{\Psi}/T_{\mathrm{max}}}$. [^7] The maximum temperature during reheating, $T_{\mathrm{max}}\sim(H_{\mathrm{inf}}M_{\mathrm{Pl}})^{1/4}T_{\mathrm{reh}}^{1/2}$ Giudice:2000ex , can be much higher than $T_{\mathrm{reh}}\sim(\Gamma_{\mathrm{inf}}M_{\mathrm{Pl}})^{1/2}$, which is the temperature at the end of reheating. Here, $H_{\mathrm{inf}}$ is the Hubble scale during the inflation, and $\Gamma_{\mathrm{inf}}$ is the decay rate of the inflaton. Once the temperature during reheating reaches $T_{\mathrm{max}}$, it decreases proportionally to $T\propto T_{\mathrm{max}}a^{-3/8}$, where $a$ is the scale factor. Even if the messengers achieve thermal equilibrium with the thermal bath, they would decouple during reheating, and their number density would be diluted by subsequent entropy production Giudice:2000ex . If $T_{\mathrm{max}}\simeq m_{\Psi}$, the relic density of the messengers is given by Giudice:2000ex
@@ Appendix B Evolution gauge couplings and kinetic mixing at high scales
In this appendix, we discuss the evolution of the gauge coupling constants and kinetic mixing in the dark GMSB model above the messenger scale. We highlight the qualitative impact of a large $\epsilon$ on gauge coupling unification through one-loop RG running. In the original basis of Eqs. (1), (2), the gauge interaction of the matter fields can be expressed as
@@
where the trace is taken over all chiral multiples.
(a) Negative $\epsilon$
(b) Positive $\epsilon$
Figure 19: The region of parameter space in which $g_{D}$ or $\epsilon$ exhibits a Landau pole below $Q=100M_{\mathrm{mess}}$ (gray shaded). The solid curve represents the bound for Scenario I, while the dashed curve indicates the bound for Scenario II. Above the blue and red curves, the pole appears at the GUT and Planck scales, respectively. Also shown are isocontours of the effective coupling $g_{\mathrm{eff}}=g_{D}\epsilon/\sqrt{(1-\epsilon^{2})}$.